OpenStudy (anonymous):

factorise 2k^2+22k+60 I got (2k+12)(k+5)

OpenStudy (anonymous):

2k^2+22k+60 2k^2+10k+12k+60 2k(k+5)+12(k+5) (2k+12)(k+5) You're correct :)

OpenStudy (anonymous):

that is what I thought, but my worksheet gave the answer 2(k+5)(k+6)

OpenStudy (anonymous):

2(k+5)(k+6) would give: 2(k^2+11k+30) = 2k^2 +22k +60 It's just another way of factorising the equation, the answer we both got is still correct, and so is the one your worksheet gives

OpenStudy (anonymous):

just factor 2 out of (2k+12)

OpenStudy (hoblos):

they are the same 2(k+5)(k+6)=(2k+12)(k+5) 2k+12 = 2(k+6)

OpenStudy (anonymous):

take 2 common from 2k+12, that is 2 ( k + 6)

OpenStudy (anonymous):

(2k+12)(k+5) = 2(k+6)(k+5)

OpenStudy (anonymous):

so in an exam, which answer would earn the marks?

OpenStudy (anonymous):

I guess (2k+12)(k+5) is not factored completely eh?

OpenStudy (hoblos):

2(k+6)(k+5) is better

OpenStudy (anonymous):

you should always factorize till u can't anymore

OpenStudy (anonymous):

I guess so, I would put the second one just to be safe, but in my maths exams the first one will get you the same amount of marks

OpenStudy (anonymous):

Great, thanks guys.